# Hydra: Unifying Document Retrieval and Generation in a Single Vision-Language Model

### 30 Mar 2026 v1

@@ Abstract
Visual document understanding typically requires separate retrieval and generation models, doubling memory and system complexity. We present Hydra, a dual-head approach that provides both ColBERT-style late-interaction retrieval and autoregressive generation from a single vision-language model (VLM). A single LoRA adapter, trained only for retrieval, is toggled at inference: enabling it produces multi-vector embeddings; disabling it recovers the base model’s generation quality—byte-identical outputs in 100% of 10,500 greedy and stochastic samples, with max $|\Delta\text{ANLS}|{=}0.0044$ across 15,301 samples on four VQA benchmarks (three informative; ChartQA is near-zero for both models under greedy decoding) when compared against an independent base-model pipeline. We identify three engineering requirements (attention-mode restoration, lm_head preservation, KV-cache-aware decoding) whose omission silently breaks generation despite correct weight recovery. On ViDoRe V1, Hydra (4B) is within 1 percentage point of a controlled single-head baseline in a single training run, with higher aggregate scores on V2 and V3 that are concentrated on a subset of tasks; multi-seed experiments are needed to confirm these trends. The single-model design reduces peak GPU memory by 41%, though adapter switching introduces throughput overhead under concurrent serving loads (§7). An ablation shows that GritLM-style joint training provides no benefit within the LoRA-based (r=16r{=}16) training regime. A proof-of-concept extension to Qwen2.5-Omni-3B demonstrates that the mechanism generalizes to audio retrieval and video embedding, with speech generation.
@@ 1 Introduction
    Evaluation on 9 ViDoRe V1 tasks against a controlled baseline, with additional single-run results on V2 (4 tasks) and V3 (8 tasks), generation equivalence across four VQA benchmarks, and efficiency measurements demonstrating 41% memory reduction (Section˜5).
@@ 3.4 Three Engineering Requirements for Dual-Head Generation
The lm_head used for generation must be the original base model’s lm_head, loaded separately from the pretrained checkpoint. Although LoRA leaves $W_{\text{base}}$ frozen in principle, in practice the lm_head can be corrupted during training through two mechanisms we identified empirically. First, when lm_head shares tied weights with the input embedding layer [Press and Wolf, 2017], gradients from the embedding propagate to lm_head even though it is not a LoRA target. Second, failing to set requires_grad=False on lm_head allows PyTorch DDP to accumulate and synchronize gradients for it even when no optimizer group updates it, causing bf16 numerical drift over thousands of steps. We avoid both failure modes by loading the lm_head from a separate instantiation of the base model and storing it alongside the adapter checkpoint.
@@ 5.4 Efficiency
Hydra (4B) uses 10,496 MB peak GPU memory during a full embed-then-generate cycle. Loading separate retrieval and generation models (ColQwen3.5 + Qwen3.5) and performing the same operations requires 17,913 MB. Hydra thus reduces peak memory by 41%.
@@ 5.5 Summary
Peak VRAM 10,496 MB (vs. 17,913 MB two-model)
@@ 6.1 Proof of Concept: Omni-Modal Generalization
We use vidore/colqwen-omni-v0.1999https://huggingface.co/vidore/colqwen-omni-v0.1, a ColBERT adapter trained on 127K image-text pairs atop the Qwen2.5-Omni-3B backbone using colpali-engine [Faysse et al., 2025]. The adapter was trained on image data only; audio and video retrieval capabilities are entirely zero-shot, acquired through the frozen Whisper audio encoder and Qwen2-VL vision encoder in the base model. We apply the Hydra architecture as-is: LoRA-on with bidirectional attention for retrieval (via custom_text_proj, 128-dim embeddings), LoRA-off with causal attention for generation (via the base model’s lm_head). No additional training is performed.
@@ 7 Discussion
Peak VRAM (single model) 10.5 GB $\sim$}10.5 GB $\times$}2 single pass full model
@@
Several directions are promising: (1) testing on non-Qwen VLM families (InternVL [Chen and others, 2024], LLaVA [Liu et al., 2023]); (2) multi-page cross-attention for document-level reasoning; (3) explicit audio and video contrastive training to improve zero-shot retrieval performance; (4) adapter composition for additional tasks beyond retrieval and generation.
@@ 8 Conclusion
Hydra demonstrates that a single retrieval-trained LoRA adapter suffices to provide both ColBERT-style document retrieval and autoregressive generation from one VLM instance, with no generation training. The key practical insight is not the toggling mechanism—which exists in prior work—but that standard training pipelines silently corrupt the base model’s generation capability through weight-tying gradients and DDP synchronization artifacts (Section˜3.4). Once these failure modes are addressed, the dual-head design matches a controlled single-head retrieval baseline within noise while preserving generation quality and reducing peak GPU memory by 41%.

### 15 Apr 2026 v2

@@ Abstract
Visual document understanding typically requires separate retrieval and generation models, doubling memory and system complexity. We present Hydra, a dual-head approach that provides both ColBERT-style late-interaction retrieval and autoregressive generation from a single vision-language model (VLM). A single LoRA adapter, trained only for retrieval, is toggled at inference: enabling it produces multi-vector embeddings; disabling it recovers the base model’s generation quality—byte-identical outputs in 100% of 10,500 greedy and stochastic samples, with max $|\Delta\text{ANLS}|{=}0.0044$ across 15,301 samples on four VQA benchmarks (three informative; ChartQA is near-zero for both models under greedy decoding) when compared against an independent base-model pipeline. We identify three engineering requirements (attention-mode restoration, lm_head preservation, KV-cache-aware decoding) whose omission silently breaks generation despite correct weight recovery. On ViDoRe V1, Hydra (4B) is within 1 percentage point of a controlled single-head baseline in a single training run, with higher aggregate scores on V2 and V3 that are concentrated on a subset of tasks; multi-seed experiments are needed to confirm these trends. The single-model design cuts peak GPU memory from 17.9 GB (two-model baseline) to 9.2 GB—a 48% reduction—though adapter switching introduces throughput overhead under concurrent serving loads (§7). An ablation shows that GritLM-style joint training provides no benefit within the LoRA-based (r=16r{=}16) training regime. A proof-of-concept extension to Qwen2.5-Omni-3B demonstrates that the mechanism generalizes to audio retrieval and video embedding, with speech generation.
@@ 1 Introduction
    Evaluation on 9 ViDoRe V1 tasks against a controlled baseline, with additional single-run results on V2 (4 tasks) and V3 (8 tasks), generation equivalence across four VQA benchmarks, and efficiency measurements showing Hydra’s 9.2 GB peak GPU memory vs 17.9 GB for an equivalent two-model deployment (Section˜5).
@@ 3.4 Three Engineering Requirements for Dual-Head Generation
The lm_head used for generation must be the original base model’s lm_head, loaded separately from the pretrained checkpoint. Although LoRA leaves $W_{\text{base}}$ frozen in principle, in practice the lm_head can be corrupted during training through two mechanisms we identified empirically. First, when lm_head shares tied weights with the input embedding layer [Press and Wolf, 2017], gradients from the embedding propagate to lm_head even though it is not a LoRA target. Second, failing to set requires_grad=False on lm_head allows PyTorch DDP to accumulate and synchronize gradients for it even when no optimizer group updates it, causing bf16 numerical drift over thousands of steps. Our specific setup sidesteps both failure modes structurally: ColQwen3.5 inherits from Qwen3_5Model (not Qwen3_5ForConditionalGeneration), so the retrieval-trained model carries no lm_head tensor at all, and the token embedding table embed_tokens is neither matched by the LoRA adapter’s target_modules nor listed in modules_to_save, so it remains byte-identical to the vanilla base through training. Since Qwen3.5 declares lm_head.weight as tied to embed_tokens.weight [Press and Wolf, 2017] via its _tied_weights_keys mapping, the weight matrix needed for generation is already resident inside the retrieval model: a thin nn.Linear instantiated on PyTorch’s meta device and then aliased to embed_tokens.weight provides the generation head at no additional memory cost, with outputs bitwise-identical (verified over 50+ greedy samples and 10 long-sequence decodes at 1024 new tokens) to loading a separate base-model instance to extract lm_head.
@@ 5.4 Efficiency
Peak GPU memory drops from 17,913 MB for an equivalent two-model deployment (ColQwen3.5 + Qwen3.5, same operations) to ${\sim}9{,}229$ MB for Hydra (4B) during a full embed-then-generate cycle—a 48% reduction. [^9] The ${\sim}1{,}267$ MB reduction vs the earlier 10,496 MB figure (a 41% reduction, pre-patch) comes from aliasing embed_tokens.weight as the generation-head lm_head.weight via Qwen3.5’s weight-tying (Section 3.4) rather than loading a second full base model to extract it. The delta matches the lm_head tensor size exactly ($|V|{\times}d{\times}2$ bytes in bf16) and produces bitwise-identical generation.
@@ 5.5 Summary
Peak VRAM ${\sim}9{,}229$ MB (vs. 17,913 MB two-model; 48% reduction)
@@ 6.1 Proof of Concept: Omni-Modal Generalization
We use vidore/colqwen-omni-v0.1101010https://huggingface.co/vidore/colqwen-omni-v0.1, a ColBERT adapter trained on 127K image-text pairs atop the Qwen2.5-Omni-3B backbone using colpali-engine [Faysse et al., 2025]. The adapter was trained on image data only; audio and video retrieval capabilities are entirely zero-shot, acquired through the frozen Whisper audio encoder and Qwen2-VL vision encoder in the base model. We apply the Hydra architecture as-is: LoRA-on with bidirectional attention for retrieval (via custom_text_proj, 128-dim embeddings), LoRA-off with causal attention for generation (via the base model’s lm_head). No additional training is performed.
@@ 7 Discussion
Peak VRAM (single model) $\sim$}9.2 GB $\sim$}9.2 GB $\times$}2 single pass full model
@@
  • Reproducibility note: In the colpali-engine 0.3.15 release we tested against, loading ColQwen3_5 directly from the public Qwen3.5-4B checkpoint random-initialises the language-model weights because the Hub checkpoint’s model.layers.* key prefix is not stripped by the package’s checkpoint-conversion mapping. The working path is to load from the adapter directory (whose model.safetensors has correctly-prefixed keys). The released Hydra checkpoint includes this full state.
@@
Several directions are promising: (1) testing on non-Qwen VLM families (InternVL [Chen and others, 2024], LLaVA [Liu et al., 2023]); (2) multi-page cross-attention for document-level reasoning; (3) explicit audio and video contrastive training to improve zero-shot retrieval performance; (4) adapter composition for additional tasks beyond retrieval and generation; (5) collapsing multi-component visual-document RAG stacks into one local VLM, trading specialist-component quality and hosted-API access for deployment locality at a capability cost we do not characterise here.
@@ 8 Conclusion
Hydra demonstrates that a single retrieval-trained LoRA adapter suffices to provide both ColBERT-style document retrieval and autoregressive generation from one VLM instance, with no generation training. The key practical insight is not the toggling mechanism—which exists in prior work—but the identification of failure modes that silently corrupt the base model’s generation capability (weight-tying gradients, DDP synchronization artifacts) together with the observation that the same weight-tying property provides a free recovery path when the architecture leaves embed_tokens frozen (Section˜3.4). Once these failure modes are addressed, the dual-head design matches a controlled single-head retrieval baseline within noise while preserving generation quality and cutting peak GPU memory roughly in half (17.9 GB $\to$ 9.2 GB for a Qwen3.5-4B class model, a 48% reduction).